OpenStudy (anonymous):

A population of insects increases at a rate of 210+8t+.6t^2 insects per day. Find the insect population after 4 days, assuming that there are 70 insects at t=0.

OpenStudy (dumbcow):

\[\frac{dP}{dt} = 210+8t+0.6t^{2}\] integrate and add initial value \[P(t) = 70+210t+4t^{2}+0.2t^{3}\] plug in t=4

OpenStudy (anonymous):

@dumbcow Thats what I though I had been doing and it says incorrect! I tried rounding and its still incorrect. Any idea why?

OpenStudy (dumbcow):

no cant think of anything at the moment so you tried 986 or 987

OpenStudy (psymon):

What answer did you try before by chance?

OpenStudy (anonymous):

986.8, 986, 987 and they all come up incorrect! I thought I was just making simple mistakes but now I'm not sure @dumbcow @Psymon

OpenStudy (psymon):

Maybe it's cranky and wants a fraction instead of a decimal? I wouldn't think you did anything wrong. If not that, then maybe you don't even bother to integrate and just plug in t = 4 and add 70 at the end. Seems silly, but when you try and answer that all 3 of us think is correct and it's wrong, then my as well throw stuff out there o.o

OpenStudy (anonymous):

I don't think thats it I actually even typed the entire equation with t plugged in and it was still incorrect. I also tried your second idea and it didn't work! Thank you both so much, I'm not sure what else it could be